OpenStudy (lgg23):

Evaluate the Indefinite Integral of sqrt(16x-x^2) dx

OpenStudy (mertsj):

Complete the square of the integrand to get 64-(x-8)^2

OpenStudy (mertsj):

Then let u = x-8 and du = dx

OpenStudy (mertsj):

Now you have; \[\int\limits_{?}^{?}\sqrt{64-u^2}du\]

OpenStudy (mertsj):

Substitute u = 8sin(s) and du = 8 cos(s)ds